# Anosov representations, strongly convex cocompact groups and weak eigenvalue gaps

### v3

@@ 8. Strongly convex cocompact subgroups of $\mathsf{PGL}_{d}(\mathbb{R})$
(ii) $\Rightarrow$ (i). We observe that $\Gamma$ is a discrete subgroup of $\mathsf{PGL}_{d}(\mathbb{R})$. Let $(\gamma_{n})_{n\in\mathbb{N}}$ be an infinite sequence of elements of $\Gamma$ and $x_{0}\in\Omega$. We may pass to a subsequence such that $\lim_{n}\gamma_{k_{n}}x_{0}\in\partial\Omega$ exists. Since $\partial\Omega$ is strictly convex we conclude that $\lim_{n}\gamma_{k_{n}}x_{0}$ is independent of the basepoint $x_{0}$. Therefore, as in [17, Lem. 7.5] or Lemma 4.5, we conclude that $\lim_{n}\frac{\sigma_{2}}{\sigma_{1}}(\gamma_{k_{n}})=0$ and $\Gamma$ has to be $P_{1}$-divergent.
Now let $(\gamma_{n})_{n\in\mathbb{N}},(\delta_{n})_{n\in\mathbb{N}}$ be two sequences of elements of $\Gamma$ which converge to the same point $x\in\partial_{\infty}\Gamma$. We claim that the limits $\lim_{n}\gamma_{n}x_{0}$, $\lim_{n}\delta_{n}x_{0}$ exist and are equal. Note that the limits will be independent of the choice of $x_{0}$. We may write
@@
where $w_{\gamma_{n}},w_{\gamma_{n}}^{\prime},w_{\delta_{n}},w_{\delta_{n}}^{\prime}\in\mathsf{PO}(d)$. Since $\Gamma$ is $P_{1}$-divergent, there exist subsequences $(\gamma_{k_{n}})_{n\in\mathbb{N}}$, $(\delta_{s_{n}})_{n\in\mathbb{N}}$ such that $a_{1}=\lim_{n}\gamma_{k_{n}}x_{0}=\lim_{n}\Xi_{1}^{+}(\gamma_{k_{n}})$, $a_{2}=\lim_{n}\delta_{s_{n}}x_{0}=\lim_{n}\Xi_{1}^{+}(\delta_{s_{n}})$, $\lim_{n}\Xi_{1}^{-}(\gamma_{k_{n}})=a_{1}^{-}$ and $\lim_{n}\Xi_{1}^{-}(\delta_{s_{n}})=a_{2}^{-}$, where $\Xi_{1}^{+}(\gamma_{k_{n}})=[w_{\gamma_{k_{n}}}e_{1}]$ and $\Xi_{1}^{-}(\gamma_{k_{n}})=[w_{\gamma_{k_{n}}}e_{d}^{\perp}]$. Proposition 6.2 and the fact that $(\gamma_{k_{n}}\cdot\delta_{s_{n}})_{\varepsilon_{1}}\rightarrow\infty$ show
@@
so $a_{i}\in a_{i}^{-}$ for $i=1,2$. In each case, the previous calculation shows that $a_{1},a_{2}\in a_{1}^{-}$ or $a_{1},a_{2}\in a_{2}^{-}$. Without loss of generality, assume that $a_{2}\in a_{1}^{-}$, so the projective line segment $[a_{1},a_{2}]$ is contained in the projective hyperplane $a_{1}^{-}$ and $\overline{\Omega}$. Since $\Gamma$ is $P_{1}$-divergent, there exist $x_{0}^{\ast}\in\Omega^{\ast}$ such that $\lim_{n}\Xi_{1}^{-}(\gamma_{k_{n}})=\lim_{n}\gamma_{k_{n}}x_{0}^{\ast}$ and $a_{1}^{-}\in\partial\Omega^{\ast}$. Therefore, $a_{1}^{-}$ avoids $\Omega$. We conclude that $[a_{1},a_{2}]$ is contained in $\partial\Omega$ and $a_{1}=a_{2}$.
@@
The implication (i) $\Rightarrow$ (ii) follows immediately by the Svarc–Milnor lemma. Now assume that (ii) holds. By [17, Thm. 1.4] it is enough to prove that $\Gamma\xhookrightarrow{}\mathsf{PGL}_{d}(\mathbb{R})$ is $P_{1}$-Anosov. Let $x_{0}\in\mathcal{C}$. Lemma 2.8 shows that the orbit map $x_{0}\mapsto\gamma x_{0}$ is a quasi-isometric embedding of $\Gamma$ into $(\mathcal{C},d_{\Omega})$, hence $\Gamma$ is word hyperbolic. By using Lemma 2.8 we deduce that there exist constants $J,k>0$ such that for every $\gamma,\delta\in\Gamma$,
$J^{-1}(\gamma\cdot\delta)_{e}-k\leqslant\big{(}\rho(\gamma)\cdot\rho(\delta)\big{)}_{\varepsilon_{1}}\leqslant J(\gamma\cdot\delta)_{e}+k.$
@@ 9. Distribution of singular values
For $q\in\mathbb{N}$ consider the $q$-symmetric power $\textup{sym}^{q}:\mathsf{GL}_{d}(\mathbb{R})\rightarrow\mathsf{GL}(\textup{Sym}^{q}(\mathbb{R}^{d}))$. Note also that with respect to the standard Cartan decomposition we have $\sigma_{1}(\textup{sym}^{q}g)=(\sigma_{1}((g))^{q}$.
@@
(i) $\rho_{R}$ is $P_{1}$-Anosov.
(ii) $\rho_{R}(\Gamma)$ is contained in a semisimple proximal Lie subgroup of $\mathsf{SL}_{d}(\mathbb{R})$ of real rank $1$.
@@
(1) The representation $\rho_{L}\times\rho_{R}:\Gamma\rightarrow\mathsf{SL}_{m+d}(\mathbb{R})$ is $P_{1}$-Anosov.
(2) $\underset{|\gamma|_{\Gamma}\rightarrow\infty}{\lim}\frac{\sigma_{1}(\rho_{L}(\gamma))}{\sigma_{1}(\rho_{R}(\gamma))}=+\infty$.
(3) $\underset{|\gamma|_{\infty}\rightarrow\infty}{\lim}\frac{\ell_{1}(\rho_{L}(\gamma))}{\ell_{1}(\rho_{R}(\gamma))}=+\infty$.
(4) There exist $C,c>0$ such that $\big{|}\log\sigma_{1}(\rho_{L}(\gamma))-\log\sigma_{1}(\rho_{R}(\gamma))\big{|}\geqslant c\log|\gamma|_{\Gamma}-C,\ \forall\gamma\in\Gamma$.
(5) There exist $C,c>0$ such that $\big{|}\log\ell_{1}(\rho_{L}(\gamma))-\log\ell_{1}(\rho_{R}(\gamma))\big{|}\geqslant c\log|\gamma|_{\infty}-C,\ \forall\gamma\in\Gamma$ of
infinite order.
@@
Let $G$ be a $P_{1}$-proximal Lie subgroup of $\mathsf{SL}_{d}(\mathbb{R})$ of real rank $1$ with Cartan projection $\mu_{G}:G\rightarrow\mathbb{R}$. Up to conjugation, we may write $G=K_{G}\exp\big{(}\mathbb{R}_{+}X_{0}\big{)}K_{G}$, $K_{G}\subset h\mathsf{SO}(d)h^{-1}$ for some $h\in\mathsf{SL}_{d}(\mathbb{R})$ and $\exp(tX_{0})=\textup{diag}(e^{ta_{1}},\ldots,e^{ta_{k}})$ with $a_{1}>a_{2}\geq\ldots\geq a_{d-1}>a_{d}$. The sub-additivity of the Cartan projection shows that there exists $M>0$ such that
$\Big{|}\log\sigma_{1}(g)-a_{i}\mu_{G}(g)\Big{|}\leqslant M$
@@
for every $g\in G$. Since we assume that $\rho_{R}$ is $P_{1}$-Anosov or $\mu(\rho_{R}(\Gamma))$ is contained in a proximal, rank $1$ Lie subgroup $G$ of $\mathsf{SL}_{d}(\mathbb{R})$, by the previous remarks we can find $A,a>0$ such that
$\log\frac{\ell_{1}(\rho_{R}(\gamma))}{\ell_{2}(\rho_{R}(\gamma))}\geqslant a\log\ell_{1}(\rho_{R}(\gamma)),\ \log\frac{\sigma_{1}(\rho_{R}(\gamma))}{\sigma_{2}(\rho_{R}(\gamma))}\geqslant a\log\sigma_{1}(\rho_{R}(\gamma))-A$
for every $\gamma\in\Gamma$.
@@
[…]
@@
for every $\gamma\in\Gamma$. Therefore, by [22, Thm. 5.3], we obtain a $\rho$-equivariant map $\xi:\partial_{\infty}\Gamma\rightarrow\mathbb{P}(\mathbb{R}^{m+d})$ which satisfies the Cartan property. Then, since $\rho(\gamma_{0})$ is $P_{1}$-proximal, we have $\xi(\gamma_{0}^{+})=\xi_{LR}^{+}(\gamma_{0}^{+})$. The minimality of the action of $\Gamma$ on $\partial_{\infty}\Gamma$ shows that $\xi=\xi_{LR}^{+}$. Then $\xi_{LR}^{+}$ satisfies the Cartan property, $\xi_{LR}^{-}$ and $\xi_{LR}^{+}$ are transverse and $\rho$ is $P_{1}$-divergent. Theorem 1.1 shows that $\rho$ is $P_{1}$-Anosov.
@@
Now suppose that $c\leqslant 1$. We choose $n\in\mathbb{N}$ large enough and consider the symmetric powers $\textup{sym}^{n}\rho_{L},\textup{sym}^{n}\rho_{R}$ of $\rho_{L},\rho_{R}$ respectively. Then $\textup{sym}^{n}\rho_{L}$ is $P_{1}$-Anosov and $\textup{sym}^{n}\rho_{R}$ satisfies either (i) or (ii). Since $\log\sigma_{1}(\textup{sym}^{n}\rho_{R}(\gamma))=n\log\sigma_{1}(\rho_{R}(\gamma))$ for $\gamma\in\Gamma$, the representation $\textup{sym}^{n}\rho_{L}\times\textup{sym}^{n}\rho_{R}$ satisfies condition (3) for $c>1$. Therefore, the previous argument implies that the representation $\textup{sym}^{n}\rho_{L}\times\textup{sym}^{n}\rho_{R}$ is $P_{1}$-Anosov. Therefore, by estimate (10), we obtain uniform constants $R,k>0$ such that ${\big{|}\log\sigma_{1}(\rho_{L}(\gamma))-\log\sigma_{1}(\rho_{R}(\gamma))\big{|}\geqslant k|\gamma|_{\Gamma}-R}$ for every $\gamma\in\Gamma$. Again we verify that $\rho$ is $P_{1}$-Anosov.
[…]
$\Big{|}\log\sigma_{1}(\rho^{ss}_{L}(\gamma))-\log\sigma_{1}(\rho^{ss}_{R}(\gamma))\Big{|}\geqslant c\log|\gamma|_{\Gamma}-L^{\prime}$
[…]
@@
Proof of Corollary 1.4. We consider the representation $\rho=\textup{sym}^{q}\rho_{1}\times\textup{sym}^{p}\rho_{2}$. The representation $\textup{sym}^{p}\rho_{2}$ is $P_{1}$-Anosov and $\textup{sym}^{q}\rho_{1}$ satisfies either condition (i) or (ii) of Theorem 9.1. The choice of $p,q\in\mathbb{N}$ shows that the representation $\textup{sym}^{p}\rho_{2}$ cannot uniformly dominate $\textup{sym}^{q}\rho_{1}$, so $\rho$ cannot be $P_{1}$-Anosov. Then, Theorem 9.1 (3) shows that for every $n\in\mathbb{N}$ we can find an element $\gamma_{n}\in\Gamma$ with $|\gamma_{n}|_{\Gamma}>n$ and $\big{|}q\mu_{1}(\rho_{1}(\gamma_{n}))-p\mu_{1}(\rho_{2}(\gamma_{n}))\big{|}\leqslant\delta\log(\mu_{1}(\rho_{1}(\gamma_{n})))$. The conclusion follows.
@@
(i) In Theorem 9.1, in the particular case where both $\rho_{L}(\Gamma)$ and $\rho_{R}(\Gamma)$ are contained in a proximal real rank $1$ Lie subgroup of $\mathsf{SL}_{m}(\mathbb{R})$ and $\mathsf{SL}_{d}(\mathbb{R})$ respectively, the equivalences $(1)\Leftrightarrow(2)\Leftrightarrow(3)$ are contained in [22, Thm. 1.14]. In the case where $\rho_{L}$ and $\rho_{R}$ take values in $\textup{Aut}_{{\mathbb{K}}}(b)$ (${\mathbb{K}}=\mathbb{R},\mathbb{C},\mathbb{H}$) for some bilinear form $b$ (see [22, §7] for background), the implications $(1)\Leftrightarrow(2)\Leftrightarrow(3)\Rightarrow(5)\Rightarrow(4)$ of Theorem 9.1 are contained in [22, Prop. 7.13 & Lem. 7.11 & Thm. 1.3].
@@
is the closed interval $\big{[}\textup{dil}_{-}(\rho_{1},\rho_{2}),\textup{dil}_{+}(\rho_{1},\rho_{2})\big{]}$. We may replace both $\rho_{1}$ and $\rho_{2}$ with their semisimplifications, and this fact also follows by the limit cone theorem of Benoist in [3, 4]. In the case where $\rho_{1}$ and $\rho_{2}$ are convex cocompact into a rank 1 Lie group, the previous fact also follows by [9, Thm. 2].

### v4

@@ 8. Strongly convex cocompact subgroups of $\mathsf{PGL}_{d}(\mathbb{R})$
(ii) $\Rightarrow$ (i). We observe that $\Gamma$ is a discrete subgroup of $\mathsf{PGL}_{d}(\mathbb{R})$. Let $(\gamma_{n})_{n\in\mathbb{N}}$ be an infinite sequence of elements of $\Gamma$ and $x_{0}\in\Omega$. We may pass to a subsequence such that $\lim_{n}\gamma_{k_{n}}x_{0}\in\partial\Omega$ exists. Since $\partial\Omega$ is strictly convex we conclude that $\lim_{n}\gamma_{k_{n}}x_{0}$ is independent of the basepoint $x_{0}$. Therefore, as in [19, Lem. 7.5] or Lemma 4.5, we conclude that $\lim_{n}\frac{\sigma_{2}}{\sigma_{1}}(\gamma_{k_{n}})=0$ and $\Gamma$ has to be $P_{1}$-divergent.
Now let $(\gamma_{n})_{n\in\mathbb{N}},(\delta_{n})_{n\in\mathbb{N}}$ be two sequences of elements of $\Gamma$ converging to $x\in\partial_{\infty}\Gamma$. We claim that the limits $\lim_{n}\gamma_{n}x_{0}$, $\lim_{n}\delta_{n}x_{0}$ exist and are equal. Note that the limits will be independent of the choice of $x_{0}$. We may write
@@
where $w_{\gamma_{n}},w_{\gamma_{n}}^{\prime},w_{\delta_{n}},w_{\delta_{n}}^{\prime}\in\mathsf{PO}(d)$. Since $\Gamma$ is $P_{1}$-divergent, there exist subsequences $(\gamma_{k_{n}})_{n\in\mathbb{N}}$, $(\delta_{s_{n}})_{n\in\mathbb{N}}$ such that $a_{1}=\lim_{n}\gamma_{k_{n}}x_{0}=\lim_{n}\Xi_{1}^{+}(\gamma_{k_{n}})$, $a_{2}=\lim_{n}\delta_{s_{n}}x_{0}=\lim_{n}\Xi_{1}^{+}(\delta_{s_{n}})$, $\lim_{n}\Xi_{1}^{-}(\gamma_{k_{n}})\\ =a_{1}^{-}$ and $\lim_{n}\Xi_{1}^{-}(\delta_{s_{n}})=a_{2}^{-}$, where $\Xi_{1}^{+}(\gamma_{k_{n}})=[w_{\gamma_{k_{n}}}e_{1}]$ and $\Xi_{1}^{-}(\gamma_{k_{n}})=[w_{\gamma_{k_{n}}}e_{d}^{\perp}]$. Proposition 6.2 and the fact that $(\gamma_{k_{n}}\cdot\delta_{s_{n}})_{\varepsilon_{1}}\rightarrow+\infty$ show
@@
so $a_{i}\in a_{i}^{-}$ for $i=1,2$. In each case, the previous calculation shows that $a_{1},a_{2}\in a_{1}^{-}$ or $a_{1},a_{2}\in a_{2}^{-}$. Without loss of generality, assume that $a_{2}\in a_{1}^{-}$, so the projective line segment $a_{1},a_{2}]$ is contained in the projective hyperplane $a_{1}^{-}$ and $\overline{\Omega}$. Since $\Gamma$ is $P_{1}$-divergent, there exist $x_{0}^{\ast}\in\Omega^{\ast}$ such that $\lim_{n}\Xi_{1}^{-}(\gamma_{k_{n}})=\lim_{n}\gamma_{k_{n}}x_{0}^{\ast}$ and $a_{1}^{-}\in\partial\Omega^{\ast}$. Therefore, $a_{1}^{-}$ avoids $\Omega$. We conclude that $a_{1},a_{2}]$ is contained in $\partial\Omega$ and $a_{1}=a_{2}$.
@@
The implication (i) $\Rightarrow$ (ii) follows immediately by the Svarc–Milnor lemma. Now assume that (ii) holds. By [19, Thm. 1.4] it is enough to prove that $\Gamma\xhookrightarrow{}\mathsf{PGL}_{d}(\mathbb{R})$ is $P_{1}$-Anosov. Let $x_{0}\in\mathcal{C}$. Lemma 2.8 shows that the orbit map $x_{0}\mapsto\gamma x_{0}$ is a quasi-isometric embedding of $\Gamma$ into $(\mathcal{C},d_{\Omega})$, hence $\Gamma$ is word hyperbolic. By using Lemma 2.8 we deduce that there exist constants $J,k>0$ such that for every $\gamma_{1},\gamma_{2}\in\Gamma$,
$J^{-1}(\gamma_{1}\cdot\gamma_{2})_{e}-k\leqslant\big(\rho(\gamma_{1})\cdot\rho(\gamma_{2})\big)_{\varepsilon_{1}}\leqslant J(\gamma_{1}\cdot\gamma_{2})_{e}+k.$
@@ 9. Distribution of singular values
Recall for $d\geq 2$, $(e_{1},\ldots,e_{d})$ denotes the canonical basis of $\mathbb{R}^{d}$. For $q\in\mathbb{N}$ consider
$\textup{Sym}^{q}\mathbb{R}^{d}:=\bigoplus_{k_{1}+\cdots+k_{d}=q}\mathbb{R}e_{1}^{k_{1}}e_{2}^{k_{2}}\cdots e_{d}^{k_{d}}$
the symmetric power of $\mathbb{R}^{d}$. The $q$-symmetric power $\textup{sym}^{q}:\mathsf{GL}_{d}(\mathbb{R})\rightarrow\mathsf{GL}(\textup{Sym}^{q}\mathbb{R}^{d}\big)$ is the representation defined as follows: for $g=(g_{ij})_{ij=1}^{n}\in\mathsf{GL}_{d}(\mathbb{R})$, define $\textup{sym}^{q}(g)(e_{1}^{k_{1}}\cdots e_{d}^{k_{d}}):=(ge_{1})^{k_{1}}\cdots(ge_{d})^{k_{d}}=\prod_{j=1}^{d}(\sum_{i}g_{ij}e_{i})^{k_{j}}$ for any basis vector $e_{1}^{k_{1}}\cdots e_{d}^{k_{d}}$ of $\textup{Sym}^{q}\mathbb{R}^{d}$.
Remark 9.1.
For $q\in\mathbb{N}$, note that respect to the standard Cartan decomposition of $\mathsf{GL}(\textup{Sym}^{q}\mathbb{R}^{d})$, for every $g\in\mathsf{GL}_{d}(\mathbb{R})$ we have that $\sigma_{1}(\textup{sym}^{q}g)=(\sigma_{1}((g))^{q}$, $\ell_{1}(\textup{sym}^{q}g)=\ell_{1}(g)^{q}$ and $\sigma_{2}(\textup{sym}^{2}g)=\sigma_{1}(g)^{q-1}\sigma_{2}(g)$, $\ell_{2}(\textup{sym}^{q}g)=\ell_{1}(g)^{q-1}\ell_{2}(g)$. In particular, by the characterizations of Anosov representations in terms of singular value (resp. eigenvalue) gaps [30, 9] (resp. [33]), a representation $\rho:\Gamma\rightarrow\mathsf{GL}_{d}(\mathbb{R})$ is $P_{1}$-Anosov if and only if $\textup{sym}^{q}\rho:\Gamma\rightarrow\mathsf{GL}(\textup{Sym}^{q}\mathbb{R}^{d}))$ is $P_{1}$-Anosov.
@@
  1. (i)
    $\rho_{R}$ is $P_{1}$-Anosov.
  2. (ii)
    $\rho_{R}(\Gamma)$ is contained in a semisimple proximal Lie subgroup of $\mathsf{SL}_{d}(\mathbb{R})$ of real rank $1$.
@@
  1. (1)
    The representation $\rho_{L}\times\rho_{R}:\Gamma\rightarrow\mathsf{SL}_{m+d}(\mathbb{R})$ is $P_{1}$-Anosov and $\rho_{L}$ uniformly dominates $\rho_{R}$.
  2. (2)
    $\underset{|\gamma|_{\Gamma}\rightarrow\infty}{\lim}\frac{\sigma_{1}(\rho_{L}(\gamma))}{\sigma_{1}(\rho_{R}(\gamma))}=+\infty$.
  3. (3)
    $\underset{|\gamma|_{\infty}\rightarrow\infty}{\lim}\frac{\ell_{1}(\rho_{L}(\gamma))}{\ell_{1}(\rho_{R}(\gamma))}=+\infty$.
  4. (4)
    There exist $C,c>0$ such that for every $\gamma\in\Gamma$ non-trivial,
    $\Big|\log\sigma_{1}(\rho_{L}(\gamma))-\log\sigma_{1}(\rho_{R}(\gamma))\Big|\geqslant c\log|\gamma|_{\Gamma}-C.$
  5. (5)
    There exist $C,c>0$ such that for every $\gamma\in\Gamma$ of infinite order
    $\Big|\log\ell_{1}(\rho_{L}(\gamma))-\log\ell_{1}(\rho_{R}(\gamma))\Big|\geqslant c\log|\gamma|_{\infty}-C.$
@@
Let $G$ be a $P_{1}$-proximal Lie subgroup of $\mathsf{SL}_{d}(\mathbb{R})$ of real rank $1$ with Cartan projection $\mu_{G}:G\rightarrow\mathbb{R}_{+}$. Up to conjugation by an element of $\mathsf{GL}_{d}(\mathbb{R})$, we may write $G=K_{G}\exp\big(\mathbb{R}_{+}X_{0}\big)K_{G}$, $K_{G}\subset h\mathsf{SO}(d)h^{-1}$ for some $h\in\mathsf{SL}_{d}(\mathbb{R})$ and $\exp(tX_{0})=\textup{diag}(e^{ta_{1}},\ldots,e^{ta_{k}})$ with $a_{1}>a_{2}\geq\ldots\geq a_{d-1}>a_{d}$. The sub-additivity of the Cartan projection shows that there exists $M>0$ such that
$\Big|\log\sigma_{i}(g)-a_{i}\mu_{G}(g)\Big|\leqslant M$
@@
for every $g\in G$. Since either (i) or (ii) holds true for $\rho_{R}$, we may find $A,a>0$ such that for every $\gamma\in\Gamma$,
$\log\frac{\ell_{1}(\rho_{R}(\gamma))}{\ell_{2}(\rho_{R}(\gamma))}\geqslant a\log\ell_{1}(\rho_{R}(\gamma)),\ \log\frac{\sigma_{1}(\rho_{R}(\gamma))}{\sigma_{2}(\rho_{R}(\gamma))}\geqslant a\log\sigma_{1}(\rho_{R}(\gamma))-A.$
@@
$\textup{(2)}\Rightarrow\textup{(1)}$. We observe that condition (2) and estimate (13) together show that $\rho$ is $P_{1}$-divergent. Since $\xi_{L}^{+}$ satisfies the Cartan property and $\sigma_{1}(\rho_{L}(\gamma))>\sigma_{1}(\rho_{R}(\gamma))$ as $|\gamma|_{\Gamma}\rightarrow\infty$, the map $\xi_{LR}^{+}$ has the Cartan property. The maps $\xi_{LR}^{+}$ and $\xi_{LR}^{-}$ are transverse, hence Theorem 1.1 shows that $\rho_{L}\times\rho_{R}$ is $P_{1}$-Anosov.
$\textup{(3)}\Rightarrow\textup{(1)}$. We are proving that $\textup{(3)}\Rightarrow\textup{(2)}$. Let $\rho_{L}^{ss},\rho_{R}^{ss}$ be semisimplifications of $\rho_{L},\rho_{R}$ respectively. By Proposition 2.6, it is enough to show that $\rho_{L}^{ss}\times\rho_{R}^{ss}$ is $P_{1}$-Anosov. By Theorem 2.5 there exists $C>0$ and a finite subset $F$ of $\Gamma$ such that for every $\gamma\in\Gamma$, there exists $f\in F$ such that $|\log\ell_{1}(\rho_{L}(\gamma f))-\log\sigma_{1}(\rho_{L}^{ss}(\gamma))|\leqslant C$ and $|\log\ell_{1}(\rho_{R}(\gamma f))-\log\sigma_{1}(\rho_{R}^{ss}(\gamma))|\leqslant C$.
Let $(\gamma_{n})_{n\in\mathbb{N}}$ be an infinite sequence of elements of $\Gamma$. For every $n$ we choose $f_{n}\in F$ satisfying the previous bounds. The triangle inequality shows $||\lambda(\rho_{L}(\gamma_{n}f_{n})||\geqslant||\mu(\rho_{L}(\gamma_{n}))||-C$, hence $\lim_{n}|\gamma_{n}f_{n}|_{\infty}=+\infty$. Therefore, $\lim_{n}\big(\log\ell_{1}(\rho_{L}^{ss}(\gamma_{n}f_{n}))-\log\ell_{1}(\rho_{R}^{ss}(\gamma_{n}f_{n}))\big)=+\infty$ so $\lim_{n}\big(\log\sigma_{1}(\rho^{ss}_{L}(\gamma_{n}))-\log\sigma_{1}(\rho_{R}^{ss}(\gamma_{n}))\big)=+\infty$. The claim now follows by $\textup{(2)}\Rightarrow\textup{(1)}$.
$\textup{(4)}\Rightarrow\textup{(1)}$. We first assume that $c>1$. By estimate (13), there exists a constant $C_{1}>0$ such that
@@
for every $\gamma\in\Gamma$. Therefore, by [24, Thm. 5.3], we obtain a $\rho$-equivariant map $\xi:\partial_{\infty}\Gamma\rightarrow\mathbb{P}(\mathbb{R}^{m+d})$ which satisfies the Cartan property. Then, since $\rho(\gamma_{0})$ is $P_{1}$-proximal, we have $\xi(\gamma_{0}^{+})=\xi_{LR}^{+}(\gamma_{0}^{+})$. The minimality of the action of $\Gamma$ on $\partial_{\infty}\Gamma$ shows that $\xi=\xi_{LR}^{+}$. Then $\xi_{LR}^{+}$ satisfies the Cartan property, $\xi_{LR}^{-}$ and $\xi_{LR}^{+}$ are transverse and $\rho$ is $P_{1}$-divergent. Theorem 1.1 shows that $\rho$ is $P_{1}$-Anosov.
@@
Now suppose $c\leqslant 1$. We choose $n\in\mathbb{N}$ large enough and consider the symmetric powers $\textup{sym}^{n}\rho_{L},\textup{sym}^{n}\rho_{R}$ of $\rho_{L},\rho_{R}$ respectively. Then $\textup{sym}^{n}\rho_{L}$ is $P_{1}$-Anosov and $\textup{sym}^{n}\rho_{R}$ satisfies either (i) or (ii). Since $\log\sigma_{1}(\textup{sym}^{n}\rho_{R}(\gamma))=n\log\sigma_{1}(\rho_{R}(\gamma))$ for $\gamma\in\Gamma$, the representation $\textup{sym}^{n}\rho_{L}\times\textup{sym}^{n}\rho_{R}$ satisfies condition (3) for $c>1$. Therefore, the previous argument implies that the representation $\textup{sym}^{n}\rho_{L}\times\textup{sym}^{n}\rho_{R}$ is $P_{1}$-Anosov. Therefore, by estimate (12), we obtain constants $R,k>0$ with
$\Big|\log\sigma_{1}(\rho_{L}(\gamma))-\log\sigma_{1}(\rho_{R}(\gamma))\Big|\geqslant k|\gamma|_{\Gamma}-R\geq 2\log|\gamma|_{\Gamma}-R$}
for all but finitely many $\gamma\in\Gamma$. Again, by the argument of the previous paaragraph, we verify that $\rho$ is $P_{1}$-Anosov.
$\textup{(5)}\Rightarrow\textup{(1)}$. It is enough to prove that the semisimplification $\rho_{L}^{ss}\times\rho_{R}^{ss}$ of $\rho$ is $P_{1}$-Anosov. Note that the representation $\rho_{L}^{ss}$ is $P_{1}$-Anosov and $\rho_{R}^{ss}$ satisfies either (i) or (ii). By Theorem 2.5 there exists $L>0$ and a finite subset $F$ of $\Gamma$ such that for every $\gamma\in\Gamma$ there exists $w\in F$ with $\lambda(\rho_{L}(\gamma w))-\mu(\rho_{L}^{ss}(\gamma))||\leqslant L$ and $\lambda(\rho_{R}(\gamma w))-\mu(\rho_{R}^{ss}(\gamma))||\leqslant L$. Since $\rho_{L}$ is a quasi-isometric embedding, by using the previous inequality, we may find $M>0$ such that $|\gamma w|_{\infty}\geqslant\frac{1}{M}|\gamma|_{\Gamma}-M$, where $\gamma\in\Gamma$ and $w\in F$ are as previously. Finally, we obtain $L^{\prime},c>0$ such that for every $\gamma\in\Gamma$ non-trivial we have
$\Big|\log\sigma_{1}(\rho^{ss}_{L}(\gamma))-\log\sigma_{1}(\rho^{ss}_{R}(\gamma))\Big|\geqslant c\log|\gamma|_{\Gamma}-L^{\prime}.$
Therefore, $\rho_{L}^{ss}\times\rho_{R}^{ss}$ is $P_{1}$-Anosov from $\textup{(4)}\Rightarrow\textup{(1)}$.
$\textup{(1)}\Rightarrow\textup{(2),(3),(4),(5)}$. Since $\ell_{1}(\rho_{L}(\gamma_{0}))>\ell_{1}(\rho_{R}(\gamma_{0}))$, $\xi_{LR}^{+}(\gamma_{0}^{+})$ is the attracting fixed point of $\rho(\gamma_{0})$ in $\mathbb{P}(\mathbb{R}^{m+d})$. The action of $\Gamma$ on $\partial_{\infty}\Gamma$ is minimal, hence $\xi_{LR}^{+}$ has to be the Anosov limit map of $\rho$ in $\mathbb{P}(\mathbb{R}^{m+d})$. In particular, $\xi_{LR}^{+}$ satisfies the Cartan property. This shows that for any sequence $(\gamma_{n})_{n\in\mathbb{N}}$ of elements of $\Gamma$ we have $\lim_{n}\big(\log\sigma_{1}(\rho_{L}(\gamma_{n}))-\log\sigma_{1}(\rho_{R}(\gamma_{n}))\big)=+\infty$. In particular, there exists $\varepsilon>0$ such that $(1-\varepsilon)\log\ell_{1}(\rho_{L}(\gamma))\geqslant\log\ell_{1}(\rho_{R}(\gamma))$ for every $\gamma\in\Gamma$. By estimates (12), (13) and Theorem 2.3 (ii) we deduce that (3), (4), (5) hold. ∎
@@
Proof of Corollary 1.4. Given $p,q\in\mathbb{N}$ with $\textup{dil}_{-}(\rho_{1},\rho_{2})\leq\frac{p}{q}\leq\textup{dil}_{+}(\rho_{1},\rho_{2})$, consider the representation $\rho_{p,q}:=\textup{sym}^{q}\rho_{1}\times\textup{sym}^{p}\rho_{2}$. The representation $\textup{sym}^{p}\rho_{2}$ is $P_{1}$-Anosov and $\textup{sym}^{q}\rho_{1}$ satisfies either condition (i) or (ii) of Theorem 9.2. The choice of $p,q\in\mathbb{N}$ shows that the representation $\textup{sym}^{p}\rho_{2}$ cannot uniformly dominate $\textup{sym}^{q}\rho_{1}$, so $\rho_{p,q}$ cannot be $P_{1}$-Anosov. Then, Theorem 9.2 (3) shows that for given $\epsilon>0$ and every $n\in\mathbb{N}$, we can find an element $\gamma_{n}\in\Gamma$ with $|\gamma_{n}|_{\Gamma}>n$ and $\big|q\mu_{1}(\rho_{1}(\gamma_{n}))-p\mu_{1}(\rho_{2}(\gamma_{n}))\big|\leqslant\epsilon\log(\mu_{1}(\rho_{1}(\gamma_{n})))$. The conclusion follows.
@@
(i) In Theorem 9.2, in the particular case where both $\rho_{L}(\Gamma)$ and $\rho_{R}(\Gamma)$ are contained in a proximal real rank $1$ Lie subgroup of $\mathsf{SL}_{m}(\mathbb{R})$ and $\mathsf{SL}_{d}(\mathbb{R})$ respectively, the equivalences $(1)\Leftrightarrow(2)\Leftrightarrow(3)$ are contained in [24, Thm. 1.14]. In the case where $\rho_{L}$ and $\rho_{R}$ take values in $\textup{Aut}_{{\mathbb{K}}}(\mathsf{B})$ ($\mathbb{K}$$\mathbb{R}$$\mathbb{C},\mathbb{H}$) for some bilinear form $\mathsf{B}$ (see [24, §7] for background), the implications $(1)\Leftrightarrow(2)\Leftrightarrow(3)\Rightarrow(5)\Rightarrow(4)$ of Theorem 9.2 are contained in [24, Prop. 7.13 & Lem. 7.11 & Thm. 1.3].
@@
is the closed interval $\big[\textup{dil}_{-}(\rho_{1},\rho_{2}),\textup{dil}_{+}(\rho_{1},\rho_{2})\big]$. We may replace both $\rho_{1}$ and $\rho_{2}$ with their semisimplifications, and this fact also follows by the limit cone theorem of Benoist in [4, 5]. In the case where $\rho_{1}$ and $\rho_{2}$ are convex cocompact into a rank 1 Lie group, the previous fact also follows by [11, Thm. 2].